TFY4340 Mesoscopic Physics Spring 2010

Updated: May 7, 2010.

Brief course summary.

• 11.01: First meeting with (some of) the students.

• 12.01: Introduction. Covered by Chapter 1 in Heinzel or Datta.

• 19.01, 25.01, 26.01, 01.02: Solid State Physics update, including the band structure of graphene. Partly covered by Chapter 2 in Heinzel, and most introductory textbooks on Solid State Physics.

• 08.02, 15.02: Surfaces, Interfaces. Covered by Chapter 3 in Heinzel.

• 22.02, 01.03: 2DEG in GaAs/AlGaAs HEMT. Covered by Chapter 3 in Heinzel.

• 01.03: Transport theory. Drude model. Boltzmann equation. Chapter 2 in Heinzel.

• 08.03: Einstein relations. Landauer formula. Chapter 7 in Heinzel, Chapter 2 in Datta.

• 15.03: Quantized conductance. (Heinzel Ch 7, Datta Ch 2) Magnetotransport. (Heinzel Ch 6, Datta Ch 1)

• 22.03: Büttiker–Landauer formalism. (Heinzel Ch 7, Datta Ch 2) Quantum Hall Effect, Edge states, Shubnikov de Haas oscillations. (Heinzel Ch 6, Datta Ch 4)

• 12.04: Quantum Interference Effects: Aharonov–Bohm effect, Weak localization, Universal conductance fluctuations. (Heinzel Ch 8, Datta Ch 3+5)

• 19.04: Single Electron Tunneling, Single electron transistor. (Heinzel Ch 9, Datta Ch 6)

• 26.04: Spintronics. Tunneling magnetoresistance (TMR), Giant magnetoresistance (GMR). (Heinzel Ch 12). (Spintronics will not be a topic on the exam in 2010.)

Exercises.

• Exercise 1, 25.01: Crystal structure. Reciprocal lattice. Brillouin zone. Density of states.

• Exercise 2, 02.02: Band structure. Tight–binding model. Graphene. 2D and 3D plotting with Matlab. Paper: P. R. Wallace, Phys Rev 71, 622 (1947).

• Exercise 3, 09.02: Effective mass. Properties of graphene.

• Exercise 4, 16.02: Surface states.

• Exercise 5, 23.02: Linear potential at AlGaAs–GaAs interface. Variational and numerical approach to the Schrödinger equation. (Matlab.)

• Exercise 6, 02.03: Boltzmann equation vs Drude model.

• Exercise 7, 16.03: 2DEG in uniform magnetic field and harmonic confining potential. Landau levels.

• Exercise 8, 13.04: 4–terminal device. Büttiker–Landauer formalism. Onsager symmetry relations. Hall resistance.

• Exercise 9, 20.04: 4–terminal device with narrow constriction. Edge states. Longitudinal resistance. Paper: H. van Houten et al, Phys Rev B 37, 8534 (1988).

• Exercise 10, 20.04: Single electron transistor.

Other relevant papers.

• B. J. van Wees et al, Phys Rev Lett 60, 848 (1988).

• D. A. Wharam et al, J Phys C 21, L209 (1988).

• K. von Klitzing et al, Phys Rev Lett 45, 494 (1980).

• Chambers, Phys Rev Lett 5, 3 (1960).

• Tonomura et al, Phys Rev Lett 48, 1443 (1982).

• Webb et al, Phys Rev Lett 54, 2696 (1985).

• Timp et al, Phys Rev Lett 58, 2814 (1987).

• McCann et al, Phys Rev Lett 97, 146805 (2006).

• Gorbachev et al, Phys Rev Lett 98, 176805 (2007).

• Skocpol et al, Phys Rev Lett 56, 2865 (1986).

• P. A. Lee, Physica 140A, 169 (1986).

• Thornton et al, Phys Rev B 36, 4514 (1987).

• Giæver and Zeller, Phys Rev Lett 20, 1504 (1968).

Department of physics, May 7, 2010. Jon Andreas Støvneng